### v1

# Utilizing Large Language Models for Information Extraction from Real Estate Transactions

Abstract
Real estate sales contracts contain crucial information for property transactions, but manual extraction of data can be time-consuming and error-prone. This paper explores the application of large language models, specifically transformer-based architectures, for automated information extraction from real estate contracts. We discuss challenges, techniques, and future directions in leveraging these models to improve efficiency and accuracy in real estate contract analysis.
1 Introduction
Real estate transactions involve complex legal documents, such as sales contracts, that outline terms and conditions agreed upon by buyers and sellers. Extracting key information from these contracts is essential for various purposes, including due diligence, risk assessment, and compliance. However, manual review and extraction of data from these documents can be labor-intensive and prone to errors.
Real estate transactions are uniquely complex due to several distinctive factors. One key aspect is the presence of contingencies, which are conditions that must be met for the transaction to proceed, such as financing or inspections. These contingencies add negotiation layers and uncertainty. Additionally, real estate transactions often involve an executory period spanning weeks or months, allowing time for inspections and repairs before final closing. Property ownership is transferred through a deed, a legal document that conveys ownership rights and must be carefully drafted and executed. Furthermore, transactions entail various liabilities like environmental issues or property defects, requiring disclosure and mitigation to minimize risk. These factors highlight the specialized expertise needed to navigate real estate transactions successfully.
In this paper, we investigate the use of large language models for extracting structured data from real estate sales contracts. Traditionally, LSTM recurrent networksSchmidhuber et al. (1997) and TransformersVaswani et al. (2017) have been widely utilized to analyze sequential data in a variety of domains. We explore techniques to preprocess documents, fine-tune models, and information extraction techniques.
2 Related Work
Recent advances in natural language processing (NLP) Chang et al. (2024), particularly with language models like BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018) and GPT (Generative Pre-trained Transformer) Brown et al. (2020), offer promising solutions for automating the extraction of information from text. These models, which have been proven to excel in a number of applications requiring sophisticated context understanding and reasoningGao et al. (2024), excel in understanding and generating human-like text, making them suitable for complex document analysis tasks.
Traditional statistical and machine learning approaches such as Conditional Random Fields (CRF), Support Vector Machines (SVM) and Hidden Markov Model (HMM) have been explored to extract named entities from legal contracts Nadeau and Sekine (2007) Betts and Jaep (2016) Surden (2021) Cui et al. (2023) Zamani and Schwartz (2017). Joshi et al. applied a sequence of information retrieval and traditional machine learning methods to determine the governing law of a contract.
While there has been ample research on large language models, the applications of machine learning models to law are scarce. To the best knowledge of the author, the applications of large language models in the domain of real estate transactions have not been well enough studied.
3 Motivations
The motivation for employing Legal Language Models (LLMs) in reading and understanding real estate transaction contracts is multifaceted and impactful. One key benefit is the optimization of attorney time. LLMs can swiftly analyze lengthy contracts, identify critical clauses, and flag potential issues, enabling attorneys to focus their efforts on higher-level legal analysis and strategic decision-making. This streamlined approach not only enhances productivity but also ensures that legal professionals can devote more time to addressing complex aspects of the transaction, ultimately providing greater value to their clients.
Furthermore, LLMs offer valuable tools for real estate agents, buyers, and sellers to comprehend contract terms effectively. By translating legal jargon into layman’s terms, LLMs empower individuals without legal expertise to understand the key provisions and implications of the contract. This enhanced understanding fosters transparency and facilitates informed decision-making during the negotiation and execution of real estate transactions. Real estate agents can better advise their clients, and buyers/sellers can navigate contractual terms with confidence, leading to smoother and more successful transactions overall.
When combined with inspection reports, appraisal reports, and past transaction history from recorded conveyance records, Large Language Models (LLMs) have the potential to consolidate a wealth of information and streamline the transaction reporting process. This integration allows LLMs to distill key insights and generate concise reports summarizing critical details such as property condition, valuation, ownership history, and legal implications. The ability of LLMs to synthesize disparate information into a cohesive and easily digestible format not only saves time but also enhances decision-making by providing stakeholders with a clear overview of the transaction’s key aspects and potential implications. Ultimately, leveraging LLMs alongside other pertinent reports and records enables a more efficient and informed approach to real estate transactions.
4 Methodology
To extract information from real estate contracts using large language models, we adopt the following methodology:
4.1 Data Preprocessing
Data processing for real estate transaction contracts within the context of large language models involves several intricate steps leveraging mathematical operations and matrix representations. Initially, raw text representing real estate contracts is tokenized into a sequence of tokens $x_{1},x_{2},\dots,x_{n}$, where each token corresponds to a word or subword unit within the document. These tokens are then mapped to unique indices using a vocabulary mapping function $\text{vocab}(x_{i})=i$, forming a tokenized input sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{n})$. Subsequently, tokens are embedded into dense vector representations using an embedding matrix $\mathbf{E}$, where each token $x_{i}$ is represented as a vector $\mathbf{e}_{i}=\mathbf{E}[i,:]$. The embedded sequence $\mathbf{X}=(\mathbf{e}_{1},\mathbf{e}_{2},\dots,\mathbf{e}_{n})$ captures semantic and contextual information crucial for understanding real estate contract text.
Additionally, positional encodings Wu et al. (2021) Vaswani et al. (2017) can be incorporated to convey sequential information in the input. This involves computing sinusoidal positional encodings based on the position $pos$ and dimension $d$ of each token, defined as:
$\text{PE}(pos,2i)=\sin\left(\frac{pos}{k^{2i/d}}\right)$
$\text{PE}(pos,2i+1)=\cos\left(\frac{pos}{k^{2i/d}}\right)$
where $i$ denotes the dimension index within the embedding vector and $k$ denotes a tuning parameter controlling the sequential relationships between texts. It’s worth noting that real estate transactions exhibit less sequential interdependence, where each paragraph generally aligns with a specific clause in a contract or a particular point in a report, resulting in relatively independent paragraphs. These positional encodings complement token embeddings and aid in capturing sequential relationships within the contract text.
Furthermore, preprocessing steps such as padding, truncation, or batch formation can ensure uniform input dimensions and facilitate efficient training and inference. These mathematical operations and matrix manipulations are fundamental in transforming raw real estate transaction text into structured numerical inputs suitable for processing by large language models, enabling tasks such as contract analysis, information extraction, and legal document understanding.
4.2 Fine-tuning Large Language Models
Fine-tuning large language models (LLMs) for real estate contract text data involves various approaches that leverage transfer learning, task-specific fine-tuning, and multi-task learning to enhance model performance and enable domain-specific learning. Transfer learning involves using pre-trained LLMs, like BERTDevlin et al. (2018) or GPTAchiam et al. (2023), trained on massive datasets for general language understanding. These models are then fine-tuned on real estate contract text data to adapt their learned features and knowledge to the nuances of real estate transactions. The fine-tuning process updates model parameters to better align with the target domain, enhancing the LLM’s ability to understand real estate-specific language and concepts.
Task-specific fine-tuning is another approach where pre-trained LLMs are fine-tuned on a dataset tailored specifically for real estate contracts. This method focuses on optimizing the LLM’s performance for a particular task, such as contract summarization or clause classification. By training the model on task-specific data, it learns to extract relevant information and make domain-specific predictions, effectively enhancing its ability to process real estate contract text.
Multi-task learning is a technique that involves training an LLM on multiple related tasks simultaneously Mahabadi et al. (2021), including real estate-specific tasks like contract interpretation and clause extraction. This approach has been used in numerous applications in specific domains Chakrabarty et al. (2019) because it encourages the model to learn shared representations across tasks, enabling it to generalize better and improve performance on each individual taskHoward and Ruder (2018) Wallingford et al. (2022). Multi-task learning helps the LLM leverage common patterns and features across different real estate contract-related tasks, leading to enhanced domain-specific learning and more robust performance.
4.3 Information Extraction
Another approach involves utilizing sequence labeling models, such as conditional random fields (CRFs), to capture structured information within real estate contracts. CRFs can model dependencies between tokens in text and assign labels to sequences corresponding to different contract elements like contingency clauses or price terms. By training LLMs with annotated real estate contract data and CRFs, the model can learn to identify and extract key information effectively.
Furthermore, LLMs can leverage semantic parsing techniques to understand the semantics of real estate contract text and extract specific attributes like property details, contract conditions, and financial terms. Semantic parsing involves mapping natural language expressions to structured representations, enabling LLMs to interpret complex contract language and extract structured information like property attributes and contractual obligations.
Mathematically, these approaches involve training LLMs with labeled real estate contract data $\mathcal{D}$ and learning to predict key information categories $\mathcal{C}$ based on the contract text $x$. Let $f(x;\theta)$ represent the LLM model parameterized by $\theta$, and $\hat{y}=f(x;\theta)$ be the predicted information categories. The objective is to minimize a loss function $\mathcal{L}(\hat{y},y)$ that measures the model’s performance in predicting the correct information categories $y$ from the real estate contract text $x$.
5 Query
Once an LLM model is fine-tuned, it is capable of answering a wide range of questions related to real estate transactions. For example, queries like ”Describe the area of the property” can be effectively answered by the model, leveraging its understanding of geographic references and property descriptions. Similarly, questions about specific contract terms, such as ”What are the contingencies in the contract?” can be addressed based on the model’s training on legal language and contract structures. By fine-tuning the LLM with relevant real estate data and legal documents, the model gains the ability to interpret and respond to diverse inquiries related to property details, transaction terms, and legal provisions, providing valuable insights and information to users involved in real estate dealings. This versatility in answering questions contributes to the efficiency and effectiveness of utilizing LLMs in real estate transactions, enhancing accessibility and understanding across various aspects of property transactions and contracts.
In fact, real estate professionals can create a tailored set of questions encompassing the crucial due diligence questions for each transaction. These questions can subsequently be inputted into the model for every deal, enabling the extraction of pertinent information and presenting key details to stakeholders involved in the transaction.
6 Qualitative Analysis
We prompted the fine-tuned LLM model with sample contracts and posed the following inquiries: ”What is the address of the property?”, ”What is the deadline for the closing?”, ”Is it possible for the buyer or seller to delay the closing by a few days?” The model accurately responded to all these questions. For the first query, the model provided the address specified in the contract: 1 Main St, New York, New York. Regarding the second question, the model correctly identified the closing deadline as January 1, 2022. For the third question, the expected response was ”no” because the contract emphasized that ”Time is of the essence.” The model accurately replied with ”no” and referenced the wording ”time is of essence.” Overall, the model was correct across all the provided sample scenarios. However, it remains to test the performance of the model on large datasets.
7 Future Directions
The application of large language models in real estate contract analysis opens up several venues for future research and development:
  • Multi-lingual Support: Extending models to handle contracts in different languages to facilitate global real estate transactions.
  • Image Support: Extending models to handle inspections and appraisal reports with images.
  • Pricing Guidance: Extending the models to provide guidance on pricing.
  • Regulatory Compliance: Integrating legal compliance checks into automated contract analysis systems.
8 Conclusion
In this paper, we have discussed the use of large language models for extracting information from real estate sales contracts. By leveraging advanced NLP techniques, we can automate tedious tasks associated with contract analysis and improve efficiency in real estate transactions. Challenges such as legal complexity and ambiguity can be addressed through domain-specific fine-tuning and ensemble learning approaches. Future research directions focus on enhancing multi-lingual support, semantic understanding, and regulatory compliance in automated contract analysis systems.